OpenStudy (booklover40):

Three to what power equals 1/387420489

OpenStudy (janu16):

is tht 1 over 387420489

OpenStudy (booklover40):

Yes.

OpenStudy (janu16):

do you have chocies

OpenStudy (booklover40):

No. I solved a bit of the question, but I need help with this part.

OpenStudy (janu16):

can you post the question instead?

OpenStudy (janu16):

i was wondering like what type of question is that lol

OpenStudy (booklover40):

Okay. Hang on...

OpenStudy (booklover40):

Three to what power is equal to (-27)^-6?

OpenStudy (booklover40):

This is for my pre-algebra class.

OpenStudy (janu16):

when you solve (-27)^-6 you get 1 over 387420489 so the question is what power to 3 gets that? hmm

OpenStudy (booklover40):

ya

OpenStudy (janu16):

alright le me think lol

OpenStudy (janu16):

(-27)^(-6)=3^x (27)^(-6)•(-1)^(-6)=3^x (27)^(-6)•(1/(-1)^(6))=3^x (27)^(-6)•(1/1)=3^x (27)^(-6)=3^x (3^3)^(-6)=3^x (3)^(-6•3)=3^x

OpenStudy (booklover40):

All I know is that the exponent is going to be negative. I have a hard time with negative exponents. :(

OpenStudy (phi):

this is a problem where you do not expand to get that huge number (making the problem very hard to figure out) they want you to use "rules of exponents" (i.e. tricks!)

OpenStudy (janu16):

@booklover40 thats what i came up with

OpenStudy (booklover40):

Thanks!

OpenStudy (janu16):

yw

OpenStudy (booklover40):

It was right! Thank you so much!

OpenStudy (anonymous):

\[\frac{1}{3^n}=\frac{1}{387420489} \]n = 18